ShadowKitten21:

factor each trinomial 4x^2+16x-63

jhonyy9:

\[4x^{2} +16x -63 = 4x^{2} +4x +4x - 4*9 = \] what you see that is common in every terms ? so you can factorize it out @XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \[4x^{2} +16x -63 = 4x^{2} +4x +4x - 4*9 = \] what you see that is common in every terms ? so you can factorize it out @XioGonz \(\color{#0cbb34}{\text{End of Quote}}\) Yessir